An electron beam passes un deflected with
uniform velocity v through two parallel
plates, when a magnetic field of 0 01 T is
applied perpendicular to an electric field
between the plates produced by a p d of
100 V The separation of the plates is 5 mm
Calculate the velocity?
1
Expert's answer
2020-05-07T06:14:42-0400

Given: "U=100 V" ; "d=5mm=5.0\cdot 10^{-3}V"; "B=0.01T=10^{-2}T"

Solution: In the problem, the motion of electrons occurs in electric and magnetic fields. In this case, two forces exert on a charged particle, these are Coulomb "\vec F_c=e\cdot \vec E" and Lorentz "\vec F_l=e\cdot[\vec vх\vec B]" forces. Since an electron beam passes un deflected with uniform velocity "\vec v" these two forces compensate each other. We may choose a coordinate system so that the electric field is directed perpendicular to plates along "\hat k" ("\vec E=E\cdot \hat k" ), the magnetic field is directed along "\hat j" ("\vec B=B\cdot \hat j" ) parallel plates and the velocity of electrons is directed along parallel plates too, say "\vec v=v_x\hat i+v_y\hat j" . If we calculate the Lorentz force,

(1) "\vec F_l=e\cdot \begin{pmatrix} \hat i & \hat j & \hat k\\ v_x & v_y &0\\0& B &0 \end{pmatrix}=e\cdot(v_x B\cdot \hat k)" we will see that the Lorentz force disappears for movement along the magnetic field. Therefore, if the electrons had a velocity component "v_y\not =0" they would move in the Y direction without Lorentz force, and a Lorentz force could not compensate the electric force. Therefore, we must put "v_y=0" and the movement of electrons occurs only along X direction. Subject to compensation of forces we have

"\vec F_c=\vec F_l" The charge of partical is reduced, the directions of forces are coincide and we get

(2) "E=v_xB" or "v_x=\frac{E}{B}"

The field "E" of parallel plates condenser can be calculated by the formula

(3)"E=\frac {U}{d}=\frac{100V}{5\cdot 10^{-3}m}=2\cdot10^4 V/m"

Substituting (3) into (2) we get finally

"v_x=\frac{2\cdot 10^4V/m}{10^{-2}T}=2\cdot 10^6 m/s"

Answer: the velocity of beam is "2\cdot 10^6 m/s"